OpenStudy (cherrilyn):

Evaluate the integral by completing the square and using trig substitution. After completing the square I got (x+3)^2-3 but I don't know what to do next

OpenStudy (yuki):

you have to give us what we are supposed to integrate is it \[\int\limits[(x+3)^2-2]dx\]?

OpenStudy (anonymous):

no, its unlikely to be that ^ lol it will be something with a division. my guess is : \[1/ [ (x+3)^2 -3] \]

OpenStudy (anonymous):

or possibly the above with a sqrt on the denominator, its a standard integral

OpenStudy (cherrilyn):

\[\int\limits_{}^{} dx \div (x ^{2} + 6x + 6 )^{2}\]

OpenStudy (cherrilyn):

oh no... Sorry.. the equation is \[\int\limits_{}^{} dx / (x ^{2} + 6x + 6)^{2} \]